# HASH-RAG: Bridging Deep Hashing with Retriever for Efficient, Fine Retrieval and Augmented Generation

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) encounters efficiency challenges when scaling to massive knowledge bases while preserving contextual relevance. We propose Hash-RAG, a framework that integrates deep hashing techniques with systematic optimizations to address these limitations. Our queries directly learn binary hash codes from knowledgebase code, eliminating intermediate feature extraction steps, and significantly reducing storage and computational overhead. Building upon this hash-based efficient retrieval framework, we establish the foundation for fine-grained chunking. Consequently, we design a Prompt-Guided Chunk-to-Context (PGCC) module that leverages retrieved hash-indexed propositions and their original document segments through prompt engineering to enhance the LLM’s contextual awareness. Experimental evaluations on NQ, TriviaQA, and HotpotQA datasets demonstrate that our approach achieves a 90% reduction in retrieval time compared to conventional methods while maintaining considerate recall performance. Additionally, The proposed system outperforms retrieval/non-retrieval baselines by 1.4-4.3% in EM scores.
@@ 1 Introduction
Figure 1: Framework of Deep Supervised Hashing with Pairwise Similarity. The framework computes similarity-preserving loss by aligning pairwise relationships between hash codes and their corresponding ground truth.
Behind the effectiveness of RAG, the huge scale of the knowledge bases is to ensure the quality and professionalism of the output Izacard et al. (2023). With the rapid growth of the model scale and the amount of knowledge data, the scale of the RAG knowledgebase that is growing increasingly has become a trend Chen et al. (2021). In view of this, the efficiency and performance of knowledgebase retrieval are more important than ever.
@@
In the domain of large-scale data retrieval, Approximate Nearest Neighbor (ANN) search attracts significant attention due to its capability to substantially reduce search complexity under most scenarios Do et al. (2016); Wang et al. (2017). Among ANN techniques Luo et al. (2021), hashing methods emerged as one of the most widely adopted approaches, offering exceptional storage efficiency and rapid retrieval capabilities Cantini et al. (2021). In particular, deep hashing methods Xia et al. (2014) employ deep neural networks to learn discriminative feature representations and convert them into compact hash codes, substantially enhancing retrieval efficiency while reducing storage requirements and computational costs Lai et al. (2015). This approach marks an unprecedented breakthrough in large-scale image retrieval and significantly outperforms conventional method Chen et al. (2021).
@@ 3.1 Hash-Based Retriever
where $l$ is the fixed hash code length and $\beta$ is the scaling hyperparameter controlling approximation smoothness ($\beta\to\infty$ converges to the sign function). We set $\sigma=0.1$ and $\beta=\sqrt{\sigma\cdot step+1}$, where $step$ counts completed training steps.
@@ 3.2 Prompt-Guided Chunk-to-Context
The compression objective minimizes $\mathcal{L}_{\mathrm{IB}}=I(\widetilde{X};X)-\beta I(\widetilde{X};Y)$, where the Lagrange multiplier $\beta$ balances information retention and compression. Unlike conventional sentence/paragraph units Karpukhin et al. (2020), we adopt proposition units Min et al. (2023) that capture atomic semantic expressions. For document $Doc$, we extract $k$ interrelated propositions $X=[x_{1},\ldots,x_{k}]$, with relevance scores computed through hybrid scoring:
@@
where $d$ is the binary code dimension. We iteratively expand the Hamming radius until selecting the top $\alpha$ propositions:
@@ 4.1 Experimental Settings
Table 1: Top $k$ recall on test sets with the index size(GB) and query time(ms) of HbR and baselines. $\dagger$ Model selected is MEVI Top-100 & HNSW from the main experiments. $\ddagger$ Integration of hash with the encoder, selecting DPR, DSH model selected is hash table lookup with candidate = 1000.
@@ 4.3 Ablations
Table 5: The effectiveness of the information bottleneck theory on the filtering data compression rate and concise mutual information in the test sets of the PGCC module for NQ and HotpotQA.
@@ 5.1 Chunk-Unit & Prompt-Guided
Information Bottleneck of Proposition   To demonstrate how HASH-RAG’s chunking strategy leverages the information bottleneck to enhance text generation capabilities, we analyze content preserved through document segmentation. Experimental analysis based on Table 5 reveals a potential correlation between compression rates and the conciseness of conditional mutual information $I(\tilde{X};X|Y;Q)$, comparing exact and greedy search methods across context lengths for generator, mutual information metrics, and EM scores. We propose that applying information bottleneck principles to factual chunking generates concise intermediate answer representations with supporting evidence, outperforming alternative strategies in multi-hop queries. This indicates conventional chunking methods cannot achieve comparable information density optimization.
@@ 5.2 Training of deep hashing algorithms
Sensitivity to Parameters   Figure 5 illustrates the hashing hyperparameter $\gamma$ sensitivity on NQ dataset with 24-bit codes. Our method shows stability across a broad range (1 < $\gamma$ < 500), with Mean Average Precision (MAP) fluctuating within 0.01. It could potentially be attributed to NQ’s hierarchical semantic structure, which exhibits tolerance to hash-induced local perturbations. This parameter invariance reduces deployment optimization complexity while ensuring multi-scenario reliability.
@@ Appendix A Prompt Template
Open-domain QA for LLaMA-2-7B Please consider all relevant details in the retrieved segments and offer a concise, informative, and contextually appropriate response. If necessary, carefully review the corresponding indexed documents to support your answer with just a few words: IdxID:23 Title: Mount Everest Propositions: Mount Everest is Earth’s highest mountain above sea level. IdxID:23 Title: Mount Everest Propositions: Mount Everest is known in Tibetan as Chomolungma. IdxID:59 Title: Lhasa Tibetan Propositions: Verbs in Tibetan always come at the end of the clause. …. ID=23 Title: Mount Everest Doc: Mount Everest, known locally as Sagarmatha or Qomolangma,[note 4] is Earth’s highest mountain above sea level, located in the Mahalangur Himal sub-range of the Himalayas. The China–Nepal border …. ID=59 Title: Lhasa Tibetan Doc: In the traditional ”three-branched” classification of the Tibetic languages, the Lhasa dialect belongs to the Central Tibetan branch (the other two being Khams Tibetan and Amdo Tibetan).[4] In terms of …. Question: What is the highest mountain in the world?
@@ B.2 Units of Wikipedia
Table 7: Statistics of text units in the English Wikipedia.

### v4

@@ Abstract
Retrieval-Augmented Generation (RAG) encounters efficiency challenges when scaling to massive knowledge bases while preserving contextual relevance. We propose Hash-RAG, a framework that integrates deep hashing techniques with systematic optimizations to address these limitations. Our queries directly learn binary hash codes from knowledgebase code, eliminating intermediate feature extraction steps, and significantly reducing storage and computational overhead. Building upon this hash-based efficient retrieval framework, we establish the foundation for fine-grained chunking. Consequently, we design a Prompt-Guided Chunk-to-Context (PGCC) module that leverages retrieved hash-indexed propositions and their original document segments through prompt engineering to enhance the LLM’s contextual awareness. Experimental evaluations on NQ, TriviaQA, and HotpotQA datasets demonstrate that our approach achieves a 90% reduction in retrieval time compared to conventional methods while maintaining considerate recall performance. Additionally, the proposed system outperforms retrieval/non-retrieval baselines by 1.4-4.3% in EM scores. Code available at https://github.com/ratSquealer/HASH-RAG.
@@ 1 Introduction
Figure 1: Framework of Deep Supervised Hashing with Pairwise Similarity. The framework computes similarity-preserving loss by aligning pairwise relationships between hash codes and their corresponding ground truth, which provide a novel perspective for optimizing RAG efficiency.
Behind the effectiveness of RAG, the huge scale of the knowledge bases is to ensure the quality and professionalism of the output Izacard et al. (2023). The remarkable effectiveness of RAG systems is fundamentally rooted in the substantial scale of the knowledge bases, which critically ensures output quality and domain expertise. Concurrently, the rapid expansion of both model architectures and knowledge data has established the continuous scaling of RAG knowledge bases Chen et al. (2021). Consequently, achieving optimal efficiency-performance coordination in knowledge base retrieval has emerged as a paramount concern in modern RAG system design.
@@
In the domain of large-scale data retrieval, Approximate Nearest Neighbor (ANN) search attracts significant attention due to its capability to substantially reduce search complexity under most scenarios Do et al. (2016); Wang et al. (2017). Among ANN techniques Luo et al. (2021), hashing methods emerged as one of the most widely adopted approaches, offering exceptional storage efficiency and rapid retrieval capabilities Cantini et al. (2021). In particular, deep hashing methods Xia et al. (2014) employ deep neural networks to learn discriminative feature representations and convert them into compact hash codes, substantially enhancing retrieval efficiency while reducing storage requirements and computational costs Lai et al. (2015). Figure 1 illustrates a typical deep supervised hashing framework. This approach marks an unprecedented breakthrough in large-scale image retrieval and significantly outperforms conventional method Chen et al. (2021), which also provide a novel perspective for optimizing RAG efficiency.
@@ 3.1 Hash-Based Retriever
where $l$ is the fixed hash code length (768 bits) and $\beta$ is the scaling hyperparameter controlling approximation smoothness ($\beta\to\infty$ converges to the sign function). We set $\sigma=0.1$ and $\beta=\sqrt{\sigma\cdot step+1}$, where $step$ counts completed training steps.
@@ 3.2 Prompt-Guided Chunk-to-Context
The compression objective minimizes $\mathcal{L}_{\mathrm{IB}}=I(\widetilde{X};X)-\beta I(\widetilde{X};Y)$, where the Lagrange multiplier $\beta$ balances information retention and compression. Unlike conventional sentence/paragraph units Karpukhin et al. (2020), we adopt proposition units Min et al. (2023) that capture atomic semantic expressions. Proposition extraction occurs during the knowledge base preprocessing phase (i.e., index construction). So for document $Doc$, we extract $k$ interrelated propositions $X=[x_{1},\ldots,x_{k}]$, with relevance scores computed through hybrid scoring:
@@
where $d$ is the binary code dimension. We iteratively expand the Hamming radius until selecting the top $\alpha$ propositions, $\alpha$ denotes the approximate candidate set retrieved via Hamming distance calculations on hash codes:
@@ 4.1 Experimental Settings
Table 1: Top $k$ recall on test sets with the index size(GB) and query time(ms) of HbR and baselines. $\dagger$ Model selected is MEVI Top-100 & HNSW from the main experiments and HNSW backbone retrieval method results in a multiplicative increase in index size, and the retrieval time of MEVI is not SOTA. $\ddagger$ Integration of hash with the encoder, selecting DPR, DSH model selected is hash table lookup with candidate = 1000.
@@ 4.3 Ablations
Table 5: The effectiveness of the information bottleneck theory on the filtering data compression rate and concise mutual information in the test sets of the PGCC module for NQ and HotpotQA. The proposition-level chunking in Hash-RAG, combined with the PGCC module guiding LLMs to focus on the correspondence between propositions and documents, is theoretically optimized and experimentally validated as the optimal trade-off between efficiency and accuracy.
@@ 5.1 Chunk-Unit & Prompt-Guided
Information Bottleneck of Proposition   To demonstrate how HASH-RAG’s chunking strategy leverages the information bottleneck to enhance text generation capabilities, we analyze content preserved through document segmentation. Experimental analysis based on Table 5 reveals a potential correlation between compression rates and the conciseness of conditional mutual information $I(\tilde{X};X|Y;Q)$, comparing exact and greedy search methods across context lengths for generator, mutual information metrics, and EM scores. We propose that applying information bottleneck principles to factual chunking generates concise intermediate answer representations with supporting evidence, outperforming alternative strategies in multi-hop queries. This indicates replacing propositions with complete documents would increased noise, distracted attention, and reduced storage efficiency, thereby degrading generation accuracy.
@@ 5.2 Training of deep hashing algorithms
Sensitivity to Parameters   Figure 5 illustrates the hashing hyperparameter $\gamma$ sensitivity on NQ dataset. Our method shows stability across a broad range (1 < $\gamma$ < 500), with Mean Average Precision (MAP) fluctuating within 0.01. It could potentially be attributed to NQ’s hierarchical semantic structure, which exhibits tolerance to hash-induced local perturbations. This parameter invariance reduces deployment optimization complexity while ensuring multi-scenario reliability.
@@ Conclusion
Acknowledgments
This rescarch is supported by the Sichuan International Science and Technology Innovation Coopcration Project with ID 2024YFHZ0317,the Chengdu Scicnce and Technology Burcau Project with ID 2024-YF09-00041-SN,and the National Natural Science Foundation of China Project with ID W2433163.
@@ Appendix A Prompt Template
Open-domain QA for LLaMA-2-7B Please consider all relevant details in the retrieved segments and offer a concise, informative, and contextually appropriate response. If necessary, carefully review the corresponding indexed documents to support your answer with just a few words: IdxID:23 Title: Mount Everest Propositions: Mount Everest is Earth’s highest mountain above sea level. IdxID:23 Title: Mount Everest Propositions: Mount Everest is known in Tibetan as Chomolungma. IdxID:59 Title: Lhasa Tibetan Propositions: Verbs in Tibetan always come at the end of the clause. …. ID=23 Title: Mount Everest Doc: Mount Everest, known locally as Sagarmatha or Qomolangma,[note 4] is Earth’s highest mountain above sea level, located in the Mahalangur Himal sub-range of the Himalayas. The China–Nepal border …. Question: What is the highest mountain in the world?
@@ B.2 Units of Wikipedia
Table 7: Statistics of text units in the English Wikipedia.